OpenStudy (anonymous):

Janet has three times as many dimes as nickels and twice as many quarters as nickels. If she has $3.40 in all, how many nickels, dimes, and quarters does she have? If n represents the number of nickels Janet has, which of the following equations could be used to solve the problem?

OpenStudy (anonymous):

Let n be nickles. N($.05)+3N($.10)+2N($.25)=$3.40.

OpenStudy (anonymous):

what about, On a trip, the fare was 50¢ for each adult and 25¢ for each child. If 30 passengers paid $12.25, how many adults and children went? Which of the following equations could not be used to solve the problem?

OpenStudy (anonymous):

Let: a = number of adults c = number of children 30 people paid $12.25 with the given fair prices per passenger: a + c = 30 0.5a + 0.25c = 12.25 two unknowns, two equations. a = 30 - c 0.5(30 - c) + 0.25c = 12.25 15 - 0.5c + 0.25c = 12.25 15 - 0.25c = 12.25 - 0.25c = -2.75 c = 11 a = 30 - c a = 30 - 11 a = 19 There were 19 adults and 11 children.

OpenStudy (anonymous):

Hope all your questions are answered! ^_^